Minutes of the Tracker Meeting

February 6, 2002

 

Action Items:

  1. Silvia: report on expected costs to do thermal vacuum testing for tower acceptance.
  2. Erik: expedite the re-machining of trays, so that they can be shipped to SLAC for mounting of payload.
  3. Ossie: order 100 more SSD dummy blanks for the engineering model.
  4. Sandro: send a sample of the silicone adhesive for encapsulation to UCSC.
  5. Erik: send documented cost estimates and thermal predictions to SLAC for the higher-conductivity sidewall fibers.

 

Sandro reported that the trays shipped from SLAC will be in Pisa tomorrow to proceed with customs on Friday.The shipment went to Milan instead of the Pisa airport.In the future SLAC needs to specify the Pisa airport in the shipping papers.One box arrived later than the others, causing further delay.The delay requires that the assembly and vibration tests be moved back a week.Sandro expects to ship the stuff back on Feb 14 or 15, to arrive back at SLAC on the following Tuesday.

 

Tom said that it turns out that foreign nationals not allowed in the Albuquerque vibration test area.Ames is another possibility.Tom has a meeting at Ames today.They are looking for work, so it may be a good match.

 

Sandro said that Silvia, Francesco, and Luisella will arrive on the afternoon of Feb 19 and Sandro will follow a day later.

 

Erik asked that we plan on 3 days for the vibration test.Sandro said that it would be convenient to go to LA on Feb 28th and 29th to visit PCI.

 

Tom reported that NASA wants to require us to do thermal-vacuum testing at the tower level for acceptance testing.4 cycles would be required.Silvia is looking into the cost, which are expected to be high given the many weeks required to cycle 18 towers.

 

Erik said that the GMSI shop received the trays for rework but hasnít looked at them yet.Tom said that he needs them early next week in order to start mounting converters, bias circuits, and ladders.

 

Tom reported on progress with the Tracker web page.All tower, tray and tooling drawings from Hytec are on Cyberdocs.Additional drawings from Pisa, especially of assembly fixtures, still need to be put there.The URL is http://www-glast.slac.stanford.edu/Tracker-Hardware/default.asp.

 

Tom reported on his trip to Teledyne last week.They now have 4 MCM boards to test their tooling with (for mounting flex circuits).Most of the polyimide boards from Young Electronics have had their edge pieces bonded on and will get their corner radius machined shortly.John is waiting for aramid material with which to do the aramid boards.All pitch adapters are at SLAC.Bias circuits should arrive today and then will be shipped to Pisa.Teledyne should have the EM boards with dummy chips built by the end of the month.

 

Tom said that 485 dummy SSDs are at SLAC waiting to be shipped to Pisa.Sandro said that he has 40 HPK mechanical samples (SSD production rejects).SLAC will order 100 more pieces of the blank wafers (3 weeks delivery).

 

The carbon-carbon closeout material is at SLAC and will be shipped to Pisa soon.BJ is working at SLAC to modify the drawings for the closeout pieces according to Luiselaís drawings.He will also work on getting the drawings for the tooling for insert bonding finalized and into Cyberdocs.

 

Robert asked if the engineering model will get thermal vacuum testing.Tom said yes, but the site not yet known.†† It was decided to discuss this when the Italians collaborators are at SLAC.

 

Sandro comments on the engineering model schedule.The Plyform detailed schedule fits fairly well with the preliminary schedule that Robert sent.There was an issue with the time to do ladder assembly.Three days per tray side had been planned, to allow for thorough curing.However, Sandro asked whether the jig could be removed after one day.Tom thinks so, based on SLAC experience.In that case the schedule would fit.Sandro will send a detailed schedule to Tom & Robert.Sandro reported that the facesheet material is at Plyform, and they will start making face sheets next week.Multiple vendors are bidding for the inserts.

 

Sandro reported that they have measured the leakage current of the Mipot ladders encapsulated with dam and fill.The results are good, with negligible increase in current and no change in depletion voltage.He will send a sample of the silicone adhesive to UCSC for long-term testing.

 

Erik reported on some issues with K1100 material.PCI is concerned that the compressive strength is half that of YS90 (the currently used material).There is less tensile strength, also, but higher modulus (so there is no problem with tower stiffness).The concern is about stress around the fasteners.We need to get a sample for pullout testing.These tests were done with the YS90 and therefore are no longer valid if we switch to K1100.Also, the lead times for the K1100 material are long: as much as 7 to 8 months to get a panel.One cannot get the material as a fabric, only unidirectional tape.Also, only 5-mil ply is available, so the sidewall needs redesigned.†† Erik suggested a Mitsubishi fiber as another possibility (it has 700 W/cm conductivity instead of 1100).Analysis predicts in that case a 4.8 C drop down the sidewalls instead of 3.6 C (the K1100 prediction).Also, there would be only a $30k to $50k price hit instead of the $200k hit we take with the K1100.Tom and Robert need to get from Hytec documented cost estimates for all engineering and test as well as the delta-T estimates so that we can make a decision.

 

Hartmut reported on SSD production.A total of 2314 SSDs are on order with 1400 delivered to Pisa and 500 on the way.Japan has purchased the expected number of 1000 this year.Also, there were 600 from Japan last year and 700 from SLAC.†† HPK has continued making 500 per month but does not have a contract for more than the 2300.INFN has not yet started ordering.A tender sent out, and 40 days are needed.They expect March 15 to make a decision.